OpenStudy (anonymous):

i will give medal is it saying X=5 or what? What is the solution to the equation –14 + x = 5? A. 19 B. 5 C. –9 D. 9

OpenStudy (anonymous):

is it sayin X=5 or what?

OpenStudy (emogirlatemoooocow):

A. 19 i think

OpenStudy (anonymous):

yeah its A

OpenStudy (emogirlatemoooocow):

@Chrissylover560212 is it right or Nah ?

OpenStudy (anonymous):

And no its not saying x=5 because x=19

OpenStudy (emogirlatemoooocow):

to explian more with these kind of questions, it is always best to isolate the x, while still keeping the equation balanced, and it will produce your answer, for example, in this question, -14 + x =5 you want to move the negative 14 to the other side of the equation, how you do this is you figure out what the opposite number to the one you are trying to move is (the opposite number is always just the same number, except for opposite negative or positive, thus, in this case, the opposite of -14 is +14 (positive 14)) take the opposite number you are trying to move, and add it to both sides, to keep the equation balanced, so if the number you want to switch is a negative, add the opposite positive, and if it is a positive, adding the negative would be the same as simply subtracting it, this process would look like this: -14 + x = 5 +14 +14 x = 5 +14 and all you have to do is figure it out from there, hope this helped peace Source(s): PS- to all the other people who answered: I'm sure he appreciated it, but why blindly give him the answer when it will help him so much more to figure out how to do these problems on his own, he's better off understanding how to do something than one problem involving it that he'll probably never see again in his life

OpenStudy (anonymous):

hold on but i got -19 @makaylaross97

OpenStudy (emogirlatemoooocow):

mhm Are you sure i got 19

OpenStudy (anonymous):

you have to do inverse operation to both sides -14+x=5 -14-(-14)+x=5-(-14) x=19 plug in 19 -14+19=5 check with a calculator if you are not sure

OpenStudy (anonymous):

yes i got -19

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hey i got another problem would -14.9+-2.6=-17.5 or is it 17.5

OpenStudy (anonymous):

Its -17.5

OpenStudy (anonymous):

okay i got it right the first time thank you @makaylaross97

OpenStudy (anonymous):

you're welcome

OpenStudy (emogirlatemoooocow):

:O

OpenStudy (anonymous):

hold on LOL

OpenStudy (anonymous):

since you helped also i fanned you okay @EmogirlAtEmoooocow

OpenStudy (emogirlatemoooocow):

Okay @Chrissylover560212

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

YAY thank you both i passed my test thanks to you 2 * dancing and jumping at the same time* LOL XD